hotdougstand:

Julie needs to cut 4 pieces of yarn, each with the same length, and a piece of yarn 7.75 inches long. Let x represent the length of each of the equal pieces of yarn that Julie decides to cut. What is the equation that can be used to determine the total length of all of the yarn that she ends up cutting, y? Is the graph of the equation continuous or discrete?

hotdougstand:

a. y=7.75x+4; discrete b. y=7.75x+4; continuous c. y=4x+7.75; discrete d. y=4x+7.75; continuou

Vocaloid:

4 pieces of yarn each with x length gives us 4*x as the length of the four pieces then we add on the 7.75 piece since x = "length of a piece" it can be any positive value, not just integers with that being said what is the equation and is it discrete or continuous?

hotdougstand:

thanks!